### v1

@@ Abstract
We propose a new specification test to assess the validity of the judge leniency design. We characterize a set of sharp testable implications, which exploit all the relevant information in the observed data distribution to detect violations of the judge leniency design assumptions. The proposed sharp test is asymptotically valid and consistent and will not make discordant recommendations. When the judge’s leniency design assumptions are rejected, we propose a way to salvage the model using partial monotonicity and exclusion assumptions, under which a variant of the Local Instrumental Variable (LIV) estimand can recover the Marginal Treatment Effect. Simulation studies show our test outperforms existing non-sharp tests by significant margins. We apply our test to assess the validity of the judge leniency design using data from Stevenson (2018), and it rejects the validity for three crime categories: robbery, drug selling, and drug possession.
@@ 1 Introduction
We propose a novel sharp test to assess the validity of the judge leniency design, which has emerged as a prominent instrumental variable (IV) approach in recent years, particularly in empirical research exploring causal effects within the criminal justice system. This design has proven beneficial in investigating the impacts of various interactions with the legal system, like pretrial detentions and incarcerations, on subsequent outcomes such as recidivism rates, conviction probabilities, and employment prospects. What sets the judge leniency design apart is its distinctive feature of randomly assigning judges to different cases, with each judge handling a significant number of cases while having discretion over the final decision. The random assignment of judges enhances the credibility of this IV approach and has led to its increasing popularity among researchers (Kling, 2006; Di Tella and Schargrodsky, 2013; Aizer and Doyle Jr, 2015; Mueller-Smith, 2015). [^6] Kling (2006) exploits randomized judge assignment along with judge propensities to instrument for incarceration length, aiming to investigate the causal impact of incarceration on labor market outcomes. Importantly, the judge leniency design’s random assignment feature extends beyond the context of criminal justice, making it a valuable methodology in diverse research contexts, including medicine, patents and startups, bankruptcy protection, evictions, and access to foster care (see Doyle Jr, Graves, Gruber, and Kleiner, 2015; Farre-Mensa, Hegde, and Ljungqvist, 2020; Dobbie, Goldsmith-Pinkham, and Yang, 2017; Gross and Baron, 2022). [^7] For example, Doyle Jr, Graves, Gruber, and Kleiner (2015) employs the judge leniency design in the medical context to examine the impact of ambulance companies on patients in emergencies, relying on the pseudo-random assignment of ambulance companies to patients. Similarly, Dobbie, Goldsmith-Pinkham, and Yang (2017) uses the leniency of randomly assigned bankruptcy judges as an instrument to study the implications of Chapter 13 bankruptcy protection on future financial events.
@@
However, in addition to the random assignment, an instrumental variable must adhere to two additional crucial conditions: (i) an exclusion restriction, which means that judges’ actions should only influence the treatment and should not have any direct influence on the defendant’s future outcomes; and (ii) a monotonicity restriction, which means that judges should consistently exhibit more or less leniency. For example, if a defendant is treated (detained) by a lenient judge, she should always be treated (detained) by a less lenient judge. While the assumption of random assignment is typically satisfied in the judge leniency design, the validity of the exclusion and monotonicity assumptions can be questionable. In practice, trial decisions (treatment) are often multidimensional, including incarceration, fines, community service, sentence length and others (Johnson, 2014). These decisions impact the future outcomes. Because different judges may have varying attitudes on these decisions, the exclusion restriction can be violated if some of the decisions are unobserved or uncontrolled. Furthermore, Abrams, Bertrand, and Mullainathan (2012) and Stevenson (2018) argue there is considerable heterogeneity in how judges rank defendants when considering various types of offences. If this heterogeneity is not observed, then it is possible that judges exhibit varying levels of leniency under different circumstances, and the monotonicity assumption would be violated. Therefore, offering a statistical test to evaluate the validity of the judge leniency design becomes a highly relevant empirical question.
In this paper, we revisit the set of testable implications of the marginal treatment effect framework derived in Heckman and Vytlacil (2005), which are essentially many inequality restrictions applied to the joint distribution of observed outcomes, treatment status, and propensity scores. We show that a specific tractable subset of these inequality restrictions characterizes the sharp testable implications of the judge leniency design. These sharp testable implications possess the unique quality of exploiting all available information within the data distribution that is useful to refute the validity of the judge leniency design. We propose asymptotically valid and consistent semi-nonparametric and semiparametric tests based on these tractable testable implications.
@@
Numerous efforts have been made to test the judge leniency design in the existing literature. A common approach involves providing separate evidence for the validity of the individual assumptions made in the judge leniency design. For instance, to assess the random assignment of judges, Dobbie, Grönqvist, Niknami, Palme, and Priks (2018) examine whether a measure of judge stringency (the instrumental variable) correlates with baseline cases and family characteristics of criminal defendants. Regarding the monotonicity assumption, they test an implication that requires the first-stage estimates to be non-negative for all subsamples. Bhuller, Dahl, Løken, and Mogstad (2018) and Norris, Pecenco, and Weaver (2021) employ similar individual testing approaches. Assessing the assumptions individually is effective in empirical scenarios where researchers know which assumption to test and have prior knowledge that other assumptions hold. However, when no such prior knowledge is available, testing assumptions individually can fail to screen out violations. In fact, the three key assumptions may collectively impose certain constraints on the observable data-generating process (DGP), which could not be detected by examining only the testable implications of each assumption in isolation.
@@
Unlike individually testing each assumption, Frandsen, Lefgren, and Leslie (2023) propose a joint test for all assumptions underlying the judge leniency design. Their test leverages the property that, in the judge leniency design, the average outcome at the judge level should exhibit a smooth relationship with the propensity score (or the judge-level treatment probability). It ought to have a bounded slope, where the bounds depend on the limits of the outcome variable’s support. Although Frandsen, Lefgren, and Leslie (2023)’s testable implication has the desirable property that it assesses all the assumptions simultaneously, we show there is still relevant information in the data distribution essential for evaluating the judge leniency design’s validity, but not used in Frandsen, Lefgren, and Leslie (2023)’s testable implication. This difference is also demonstrated by the simulation and empirical studies reported in Section 4.
@@
To the best of our knowledge, our test is the only sharp test available for assessing the validity of the judge leniency design. In other words, our testable implications exhaust all the information in the observed data distribution. As seen in previous methods, non-sharp tests have practical virtue when there is no easily tractable characterization of the sharp testable implications of a model’s assumptions. If a non-sharp rejects, it conveys an informative result that the assumptions should be rejected. However, there are also important trade-offs to consider. First, a non-sharp test can have no power against certain violations since it does not consider all possible constraints on the data distribution. Second, different non-sharp tests can lead to discordant empirical results and potentially misleading interpretations of the estimand of interest (see Kédagni, Li, and Mourifié, 2020). For instance, two different non-sharp tests may produce conflicting results because they consider different aspects of the observed data distribution. Our sharp test addresses both issues as it is a consistent test built upon sharp testable implications and, therefore, a useful complement to the existing tests.
@@
As a potential alternative to the existing non-sharp tests, one may consider testing the validity of the judge leniency design employing some of the existing sharp tests developed for the Local Average Treat Effect (LATE) framework, i.e., Kitagawa (2015), Huber and Mellace (2015), and Mourifié and Wan (2017). However, it is worth noting these tests may over-reject since they are based on a priori direction in the monotonicity assumption and are not directly applicable in the context of judge leniency design. For instance, in the judge leniency design, the number of judges can be quite large, and in some cases, it might even be infinite, especially when judges’ types are continuous. In such scenarios, the number of potential directions to consider becomes large, possibly infinite. Imposing a specific ex-ante direction in the judge leniency design is therefore overly restrictive, and to consider all possible directions might be impractical or impossible. Furthermore, imposing an incorrect a priori direction bears an additional risk of model misspecification. These issues highlight the need for a more flexible testing approach, like the one proposed in this paper, which is free from making overly restrictive assumptions on the direction of monotonicity.
@@
Finally, while our test is mainly motivated by testing judge leniency designs, it can be applied to test the identifying assumptions in a general Marginal Treatment Effect framework with continuous or discrete instrument variables, which has been applied to various empirical settings. See Carneiro, Heckman, and Vytlacil (2011); Kowalski (2016); Brinch, Mogstad, and Wiswall (2017), among many others.
@@ 2 Model and Sharp Testable Implications
We adopt the potential outcomes framework. Let the observed treatment indicator be $D\in\{0,1\}$. For example, in the judge leniency design, $D=1$ denotes incarceration. We denote by $Z$ the judge’s observable type and let $Z\in\mathcal{Z}\subseteq\mathbb{R}^{d_{z}}$ for $d_{z}\geq 1$. $Y_{d}(z)\in\mathcal{Y}\subseteq\mathbb{R}$ denotes the potential outcome of interests (e.g. recidivism) when the treatment and the judge’s type are externally set to $D=d$, and $Z=z$, respectively. Similarly, $D(z)$ denotes the potential treatment when the judge’s type is externally set to $Z=z$. Let $Y=Y_{1}(Z)D+Y_{0}(Z)(1-D)$ be the observed outcome. For the moment, we omit observed defendant and case covariates $X$ (such as time and courtroom of the trial) for ease of notation. The identification analysis in this section can be extended by conditioning on $X$. We will also discuss the implementation of our test in the presence of $X$ in Section 3.2.
In our setting, $Z$ can be multidimensional; it can also be continuous, discrete, or mixed. For example, if there is a group of judges $\mathcal{J}$, and their identities are observed, then $Z\in\mathcal{J}$ can be chosen as the identity of the judge assigned to the defendant. This is the instrumental variable that Frandsen, Lefgren, and Leslie (2023, FLL hereafter) consider. On the other hand, we allow scenarios in which the judge’s identity is unobserved but with observed characteristics. In this case, $Z$ may contain a set of continuous or discrete variables, such as the judge’s experience, gender, and race. The literature mainly relies on the following assumptions to evaluate the causal effects of treatment $D$ on outcome $Y$.
@@
For any pair $(z,z^{\prime})\in\mathcal{Z}\times\mathcal{Z}$ either $D(z)\geq D(z^{\prime})$ for all defendants or $D(z)\leq D(z^{\prime})$ for all defendants.
A particular feature of the judge leniency design is judges are usually randomly assigned to different cases, making the random assignment assumption likely to hold in practice. Hence, we assume that Assumption 2.1 holds throughout the paper. However, Assumptions 2.2 and 2.3 are usually less credible. Assumption 2.2 means the effect of judges on the potential outcomes must necessarily transit through their effect on treatment assignment. Assumption 2.3 requires any defendants treated (incarcerated) by a more lenient judge be also treated if assigned to a less lenient one. Heckman and Vytlacil (2005) refer to the monotonicity assumption as a uniformity condition since it imposes a restriction across judges rather than the shape of a function for a particular judge. Vytlacil (2002) provides an equivalent characterization of the monotonicity assumption, which can be stated as follows:
@@
$Y$ $=$ $Y_{1}D+Y_{0}(1-D),$ (2.1)
$D$ $=$ $1\left\{P(Z)\geq V\right\}.$ (2.2)
Assumptions 2.1, 2.2 and 2.4 (equivalently Assumptions 2.1, 2.2 and 2.3) impose some restrictions on the joint distribution of the observed variables $(Y,D,P(Z))$, which we will characterize in Theorem 2.1. But before stating the theorem, let us give the intuition of the testable implications. Let $g:\mathcal{Y}\rightarrow\mathbb{R}^{+}$ be a nonnegative real integrable function such that $\mathbb{E}|g(Y_{d})|<\infty$. Taking $d=0$ as an illustration. For any pair $(p,p^{\prime})\in\mathcal{P}\times\mathcal{P}$ such that $p\leq p^{\prime}$, we have:
@@
(i) If Assumptions 2.1, 2.2 and 2.4 (equivalently Assumptions 2.1, 2.2 and 2.3) hold, then for all $y,y^{\prime}\in\mathcal{Y}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p)$ and $-\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p)$ are non-decreasing in $p$ for all $p\in\mathcal{P}$.
@@
(ii) If for all $y,y^{\prime}\in\mathcal{Y}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p)$ and $-\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p)$ are non-decreasing in $p$ for all $p\in\mathcal{P}$, there exists a joint distribution of $(\tilde{V},\tilde{Y}_{1},\tilde{Y}_{0},P(Z))$ such that Assumptions 2.1, 2.2 and 2.4 hold, and $(\tilde{Y},\tilde{D},P(Z))$ has the same distribution as $(Y,D,P(Z))$.
@@
The proof of Theorem 2.1 is collected in the supplementary materials Section B. The testable implications in Theorem 2.1(i) are a subset of the implications previously derived in Heckman and Vytlacil (2005, Appendix A), who show for any non-negative integrable function, i.e. $g(\cdot):\mathcal{Y}\rightarrow\mathbb{R}^{+}$, $\mathbb{E}[g(Y)D|P=p]$ and $-\mathbb{E}[g(Y)(1-D)|P=p]$ are non-decreasing in $p$ under Assumptions 2.1, 2.2 and 2.4. The contribution of Theorem 2.1-(i) is that it shows we do not need to visit every single non-negative measurable function. It is sufficient to restrict our attention to a tractable sub-class of these functions to screen all possible observable violations. This tractable characterization provides a basis for constructing a formal statistical test to verify the validity of the assumptions.
@@
The second part of Theorem 2.1 is new, and it shows the testable implications in Theorem 2.1(i) are the most informative way to detect all observable violations of the random assignment, the exclusion restriction, and the monotonicity assumption (without an ex-ante imposed direction). These testable implications cannot be strengthened without making additional assumptions. Various tests or testable implications are used in the literature to screen violations of the judge leniency design assumptions; for instance, Dobbie, Grönqvist, Niknami, Palme, and Priks (2018); Bhuller, Dahl, Løken, and Mogstad (2018); Norris, Pecenco, and Weaver (2021); Frandsen, Lefgren, and Leslie (2023). However, to the best of our knowledge, only Theorem 2.1 provides sharp testable implications without imposing an a priori direction in the monotonicity assumption.
@@
Tests based on sharp testable implications have empirical virtue. In practice, one may use tests developed from non-sharp testable implications for the sake of traceability. However, as recently discussed in Kédagni, Li, and Mourifié (2020), non-sharp tests can lead to discordant empirical results and misleading interpretations of the estimand of interest. It is possible that for the same data, two different non-sharp tests generate contradictory results as they may use different sets of information from the same observed DGP to screen violations of the model assumptions. Thus, the conclusion may largely depend on which test the empirical researcher implements.
@@
Moreover, after implementing a specification test and obtaining a non-rejection result, one often proceeds and provides a causal interpretation of the estimand. For example, in judge leniency designs, the 2SLS or Local IV (LIV) estimand is interpreted as the LATE or MTE, respectively. However, since a non-sharp test only uses part of the observable information in the data and fails to reject the model when it is misspecified, we must be cautious about interpreting the 2SLS or the Local IV estimand as identifying the LATE/MTE solely based on the result of a non-sharp test. Therefore, using a sharp test must be viewed not only as a theoretical exercise, but also as having important empirical relevance. A sharp test provides the most informative way to detect all observable violations of a given model’s assumptions and is more robust to possible misleading interpretations and discordant results.

### v2

@@ Abstract
We propose sharp testable implications and tests to jointly assess the random assignment, exclusion, and monotonicity assumptions in judge leniency designs. Our procedures accommodate various data scenarios in which the number of defendants handled by a judge may be either small or large, and allow for discrete or continuous instrumental variables. When the validity of the design is rejected, a variant of the marginal treatment effect can be identified under weaker assumptions. We apply our test to the Philadelphia court data studied by Stevenson (2018) and demonstrate that it outperforms non-sharp joint tests by significant margins in simulation studies.
@@ 1 Introduction
We propose a novel sharp test to assess the validity of the judge leniency design, which has emerged as a prominent instrumental variable (IV) approach in recent years, particularly in empirical research exploring causal effects within the criminal justice system. This design has proven beneficial in investigating the impacts of various interactions with the legal system, such as pretrial detentions and incarcerations, on subsequent outcomes, including recidivism rates, conviction probabilities, and employment prospects. What sets the judge leniency design apart is its distinctive feature of randomly assigning judges to different cases, with each judge handling a significant number of cases while having discretion over the final decision. The random assignment of judges enhances the credibility of this IV approach and has led to its increasing popularity among researchers (Kling, 2006; Di Tella and Schargrodsky, 2013; Aizer and Doyle Jr, 2015; Mueller-Smith, 2015). [^6] Kling (2006) exploits randomized judge assignment along with judge propensities to instrument for incarceration length, aiming to investigate the causal impact of incarceration on labor market outcomes. Importantly, the judge leniency design’s random assignment feature extends beyond the context of criminal justice, making it a valuable methodology in diverse research contexts, including medicine, patents and startups, bankruptcy protection, evictions, and access to foster care (see Doyle Jr, Graves, Gruber, and Kleiner, 2015; Farre-Mensa, Hegde, and Ljungqvist, 2020; Dobbie, Goldsmith-Pinkham, and Yang, 2017; Gross and Baron, 2022). [^7] For example, Doyle Jr, Graves, Gruber, and Kleiner (2015) employs the judge leniency design in the medical context to examine the impact of ambulance companies on patients in emergencies, relying on the pseudo-random assignment of ambulance companies to patients. Similarly, Dobbie, Goldsmith-Pinkham, and Yang (2017) uses the leniency of randomly assigned bankruptcy judges as an instrument to study the implications of Chapter 13 bankruptcy protection on future financial events.
@@
However, in addition to the random assignment, an instrumental variable must adhere to two additional crucial conditions: (i) an exclusion restriction, which means that judges’ actions should only influence the treatment and should not have any direct influence on the defendant’s future outcomes; and (ii) a monotonicity restriction, which means that judges should consistently exhibit more or less leniency. This means that if a defendant was treated (detained) by one judge, she would always be treated (detained) by a less lenient judge. Trial decisions (treatment) are often multidimensional, including incarceration, fines, community service, sentence length, and others (Johnson, 2014). These decisions impact future outcomes. Because different judges may have varying attitudes on these decisions, the exclusion restriction can be violated if some of the decisions are unobserved or uncontrolled. Furthermore, Abrams, Bertrand, and Mullainathan (2012) and Stevenson (2018) argue there is considerable heterogeneity in how judges rank defendants when considering various types of offenses. If this heterogeneity is not observed, then it is possible that judges exhibit varying levels of leniency under different circumstances, and the monotonicity assumption would be violated. These observations align with Mogstad, Torgovitsky, and Walters (2019), who demonstrate that, in general, monotonicity effectively requires homogeneous choice behavior for economic agents when there are multiple instruments. Therefore, offering a statistical test to evaluate the validity of the judge leniency design becomes a highly relevant empirical question.
In this paper, we characterize the sharp testable implications of the judge leniency design as a set of inequality restrictions on the distribution of the observed data. Our result is novel and contributes to the testable implications derived in the seminal work of Heckman and Vytlacil (2005) in two ways. First, our implications belong to a tractable subset of the constraints of Heckman and Vytlacil (2005) and are easier to implement in practice. Second, we establish the sharpness of our testable implication, that is, they possess the unique quality of exploiting all available information within the data distribution that is useful to refute the validity of the judge leniency design.
@@
Numerous efforts have been made to test the judge leniency design in the existing literature. A common approach involves providing separate evidence for the validity of the individual assumptions made in the judge leniency design. For instance, to assess the random assignment of judges, Dobbie, Grönqvist, Niknami, Palme, and Priks (2018) examines whether a measure of judge stringency (the instrumental variable) correlates with baseline cases and family characteristics of criminal defendants. Regarding the monotonicity assumption, they test an implication that requires the first-stage estimates to be non-negative for all subsamples. Bhuller, Dahl, Løken, and Mogstad (2018) and Norris, Pecenco, and Weaver (2021) employ similar individual testing approaches. Assessing the assumptions individually is effective in empirical scenarios where researchers know which assumption to test and have prior knowledge that other assumptions hold. Our approach adds to the existing body of knowledge by introducing a test that does not depend on prior information. In fact, the three key assumptions may collectively impose certain constraints on the observable data-generating process (DGP), which could not be detected by examining only the testable implications of each assumption in isolation.
@@
Unlike individually testing each assumption, Frandsen, Lefgren, and Leslie (2023) proposes a joint test for all assumptions underlying the judge leniency design. Their test leverages the property that, in the judge leniency design, the average outcome at the judge level should exhibit a smooth relationship with the propensity score (or the judge-level treatment probability). It ought to have a bounded slope, where the bounds depend on the limits of the outcome variable’s support. Although Frandsen, Lefgren, and Leslie (2023)’s testable implication has the desirable property that it assesses all the assumptions simultaneously, we show there is still significant relevant information in the data distribution essential for evaluating the judge leniency design’s validity, but not used in Frandsen, Lefgren, and Leslie (2023)’s testable implication. This difference is also demonstrated by numerical examples and empirical studies reported in Sections 2 and 4.
@@
To the best of our knowledge, our test is the only sharp test available for assessing the validity of the judge leniency design. In other words, our testable implications exhaust all the information in the observed data distribution. As seen in previous methods, non-sharp tests have practical virtue when there is no easily tractable characterization of the sharp testable implications of a model’s assumptions. If a non-sharp rejects, it conveys an informative result that the assumptions should be rejected. However, there are also important trade-offs to consider. First, a non-sharp test can have no power against certain violations since it does not consider all possible constraints on the data distribution. Second, different non-sharp tests can lead to discordant empirical results and potentially misleading interpretations of the estimand of interest (see Li, Kédagni, and Mourifié, 2024). For instance, two different non-sharp tests may produce conflicting results because they consider different aspects of the observed data distribution. Our sharp test addresses both issues as it is a consistent test built upon sharp testable implications and, therefore, a useful complement to the existing literature.
We construct valid and consistent semi-nonparametric and semiparametric tests based on these tractable testable implications. Our asymptotic tests support a diverse range of data structures. For example, we can apply our tests to the empirical context in which each judge handles a large number of defendants, and the number of judges can be either large or small (as in our empirical application). As we will further elaborate in Section 3, our asymptotic tests are also applicable when the number of defendants for each judge is small, as long as the data regime permits a root-n estimation of the propensity score.
We also provide an easy-to-compute finite sample test for cases involving a small number of judges and a small number of defendants per judge. To the best of our knowledge, ours and the finite sample test of Frandsen, Lefgren, and Leslie (2023) are the only finite sample specification tests in the judge leniency design literature. Like Frandsen, Lefgren, and Leslie (2023), our finite sample test also focuses on binary outcomes. Unlike Frandsen, Lefgren, and Leslie (2023)’s test, which ensures the finite sample validity by computing a “least favorable p-value” via a high-dimensional nonlinear optimization routine, we use Bonferroni correction. The computation for our test is very light and requires little more than simulating Bernoulli random variables. Therefore, it serves as a useful addition to the existing finite sample tests.
@@
As a potential alternative to the existing non-sharp tests, one may consider testing the validity of the judge leniency design employing some of the existing sharp tests developed for the Local Average Treatment Effect (LATE) framework, i.e., Kitagawa (2015), Huber and Mellace (2015), and Mourifié and Wan (2017). However, it is worth noting that these tests may over-reject since they are based on a priori direction in the monotonicity assumption and are not directly applicable in the context of judge leniency design. For instance, in the judge leniency design, the number of judges can be quite large, and in some cases, it might even be infinite, especially when judges’ types are continuous. In such scenarios, the number of potential directions to consider becomes large, possibly infinite. Imposing a specific ex-ante direction in the judge leniency design is therefore overly restrictive, and considering all possible directions might be impractical or impossible. Furthermore, imposing an incorrect a priori direction bears an additional risk of model misspecification. These issues highlight the need for a more flexible testing approach, like the one proposed in this paper, which is free from making overly restrictive assumptions on the direction of monotonicity.
@@
While our test is primarily motivated by testing judge leniency designs, it can also be applied to assess the identifying assumptions in a general Marginal Treatment Effect framework with continuous or discrete instrument variables, which has been applied to various empirical settings. See Carneiro, Heckman, and Vytlacil (2011); Kowalski (2016); Brinch, Mogstad, and Wiswall (2017), among many others. In the context of judge leniency designs, this also means that our test does not require observing a judge’s identity and accommodates continuous judge types. Finally, motivated by Mogstad, Torgovitsky, and Walters (2019), we propose to relax monotonicity and exclusion assumptions to partial monotonicity and partial exclusion, respectively, when our test rejects the null hypothesis.
@@ 2 Model and Sharp Testable Implications
We adopt the potential outcomes framework. Let the observed treatment indicator be $D\in\{0,1\}$. For example, in the judge leniency design, the unit of observation is defendants. Hence, $D=1$ indicates that a defendant is incarcerated. Let $Z\in\mathcal{Z}\subseteq\mathbb{R}^{d_{z}}$ be the type of the judge assigned to the defendant. $Y_{d}(z)\in\mathcal{Y}\subseteq\mathbb{R}$ denotes the potential outcome of interest (e.g., recidivism) when the treatment and the judge’s type are externally set to $D=d$, and $Z=z$, respectively. Similarly, $D_{z}$ denotes the potential treatment when the judge’s type is externally set to $Z=z$. Let $Y=Y_{1}(Z)D+Y_{0}(Z)(1-D)$ be the observed outcome. For the moment, we omit observed defendant and case covariates $X$ (such as time and courtroom of the trial) for ease of notation. The identification analysis in this section can be extended by conditioning on $X$. We will also discuss the implementation of our test in the presence of $X$ in Section 3.2.
In our setting, $Z$ can be multidimensional, continuous, discrete, or a combination of both. For example, if there is a group of judges $\mathcal{J}$, and if their identities are observed, then $Z\in\mathcal{J}$ can be chosen as the identity of the judge assigned to the defendant. This is the instrumental variable that Frandsen, Lefgren, and Leslie (2023, FLL hereafter) consider. On the other hand, we allow scenarios in which the judge’s identity is unobserved but with observed characteristics. In this case, $Z$ may contain a set of continuous or discrete variables, such as the judge’s experience, gender, and race.
The literature mainly relies on the following assumptions to evaluate the causal effects of treatment $D$ on outcome $Y$.
@@
For any pair $(z,z^{\prime})\in\mathcal{Z}\times\mathcal{Z}$ either $D_{z}\geq D_{z^{\prime}}$ for all defendants or $D_{z}\leq D_{z^{\prime}}$ for all defendants.
A particular feature of the judge leniency design is that judges are usually randomly assigned to different cases, making the random assignment assumption likely to hold in practice. However, Assumptions 2.2 and 2.3 are usually less credible. Assumption 2.2 means the effect of judges on the potential outcomes must necessarily transit through their effect on treatment assignment. Assumption 2.3 requires that any defendants treated (incarcerated) by a more lenient judge be also treated if assigned to a less lenient one. Heckman and Vytlacil (2005) refers to the monotonicity assumption as a uniformity condition since it restricts that the treatment on all the defendants must vary in a uniform direction when externally assigned to another judge. Vytlacil (2002) provides an equivalent characterization of the monotonicity assumption, which can be stated as follows:
@@
$\begin{cases}Y=Y_{1}D+Y_{0}(1-D),\\ D=1\left\{P(Z)\geq V\right\}.\end{cases}$ (2.1)
Assumptions 2.1, 2.2 and 2.4 (equivalently Assumptions 2.1, 2.2 and 2.3) impose some restrictions on the joint distribution of the observed variables $(Y,D,P(Z))$, which we will characterize in Theorem 2.1. But before stating the theorem, we will discuss the intuition of the testable implications. Let $g:\mathcal{Y}\rightarrow\mathbb{R}^{+}$ be a nonnegative real integrable function such that $\mathbb{E}|g(Y_{d})|<\infty$. Taking $d=0$ as an illustration. For any pair $(p,p^{\prime})\in\mathcal{P}\times\mathcal{P}$ such that $p\leq p^{\prime}$, we have:
@@
(i) If Assumptions 2.1, 2.2 and 2.4 (equivalently Assumptions 2.1, 2.2 and 2.3) hold, then for all $y,y^{\prime}\in\mathcal{Y}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p)$ and $\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p)$ are non-decreasing in $p$ for all $p\in\mathcal{P}$.
@@
(ii) If for all $y,y^{\prime}\in\mathcal{Y}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p)$ and $\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p)$ are non-decreasing in $p$ for all $p\in\mathcal{P}$, there exists a joint distribution of $(\tilde{V},\tilde{Y}_{1},\tilde{Y}_{0},P(Z))$ such that Assumptions 2.1, 2.2 and 2.4 hold, and $(\tilde{Y},\tilde{D},P(Z))$ has the same distribution as $(Y,D,P(Z))$.
@@
The proof of Theorem 2.1 is collected in Section B.1. The testable implications in Theorem 2.1(i) are a subset of the implications previously derived in Heckman and Vytlacil (2005, Appendix A), who show for any non-negative integrable function, i.e. $g(\cdot):\mathcal{Y}\rightarrow\mathbb{R}^{+}$, $\mathbb{E}[g(Y)D|P=p]$ and $\mathbb{E}[g(Y)(1-D)|P=p]$ are non-decreasing in $p$ under Assumptions 2.1, 2.2 and 2.4. The contribution of Theorem 2.1-(i) is that it shows we do not need to visit every single non-negative measurable function. It is sufficient to restrict our attention to a tractable subclass of these functions to screen all possible observable violations. This tractable characterization provides a basis for constructing a formal statistical test to verify the validity of the assumptions. [^8] We note that use the the half-interval class $g(Y)=1\{Y\leq y\},y\in\mathcal{Y}$ will result in loss of power. To see this, suppose the support is finite, that is, $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{K}\}$, then it is without loss of information to consider the class of singletons: $g(Y)=1\{Y=y_{k}\}$, $k=1,2,\cdots,K$. However, if one considers $g(Y)=1\{Y\leq y_{k}\}$, then it is possible that both $\mathbb{P}(Y\leq y_{1},D=1|P=p)$ and $\mathbb{P}(Y\leq y_{2},D=1|P=p)$ are non-decreasing function, but $\mathbb{P}(Y=y_{2},D=1|P=p)$ is not.
@@
The second part of Theorem 2.1 is new, and it shows that the testable implications in Theorem 2.1(i) are the most informative way to detect all observable violations of the random assignment, the exclusion restriction, and the monotonicity assumption (without an ex-ante imposed direction). These testable implications cannot be strengthened without making additional assumptions. Various tests or testable implications are used in the literature to screen violations of the judge leniency design assumptions; for instance, Dobbie, Grönqvist, Niknami, Palme, and Priks (2018); Bhuller, Dahl, Løken, and Mogstad (2018); Norris, Pecenco, and Weaver (2021); Frandsen, Lefgren, and Leslie (2023). However, to the best of our knowledge, only Theorem 2.1 provides sharp testable implications without imposing an a priori direction in the monotonicity assumption.
@@
Tests based on sharp testable implications have empirical virtue. In practice, one may use tests developed from non-sharp testable implications for the sake of traceability. However, as recently discussed in Li, Kédagni, and Mourifié (2024), non-sharp tests can lead to discordant empirical results and misleading interpretations of the estimand of interest. It is possible that for the same data, two different non-sharp tests may generate contradictory results, as they may use different sets of information from the same observed DGP to screen violations of the model assumptions. Thus, the conclusion may largely depend on which test the empirical researcher implements.
@@
Moreover, after implementing a specification test and obtaining a non-rejection result, one often proceeds and provides a causal interpretation of the estimand. For example, in judge leniency designs, the 2SLS or Local IV (LIV) estimand is interpreted as the LATE or MTE, respectively. However, since a non-sharp test only uses part of the observable information in the data and fails to reject the model when it is misspecified, we must be cautious about interpreting the 2SLS or the LIV estimand as identifying the LATE/MTE solely based on the result of a non-sharp test. Therefore, using a sharp test must be viewed not only as a theoretical exercise, but also as having an important empirical relevance. A sharp test provides the most informative way to detect all observable violations of a given model’s assumptions and is more robust to possible misleading interpretations and discordant results.